OpenStudy (anonymous):

DFG and JKL are complementary angles. m DFG= x+5, and m JKL=X-9. find the measure of each angle.

OpenStudy (anonymous):

complementary means that the angles add up to 90.. x+5 + x- 9 = 90 2x - 4= 90 94=2x x= 47

OpenStudy (anonymous):

now plug in x to both of the angles equations and theres your answer

OpenStudy (anonymous):

set them equal 90 degrees because complementary means they form 90 degrees. x + 5 + x - 9 = 90 2x - 4 = 90 +4 +4 ---- ----- 2x = 90 __ __ 2 2 x = 45